OpenStudy (anonymous):

I WILL FAN AND GIVE MEDALS!!!1 PLZ HELP!!!! Use the substitution method to solve the following system of linear equations: 2x + 2y = 10 x = –y + 5 A. There is no solution. B. There are an infinite number of solutions. C. The solution is (4, –1). D. The solution is (2, –3).

OpenStudy (anonymous):

I am thinking it would be (4, -1)

OpenStudy (anonymous):

put x=-y+5 from the second equation into the first, so that the result is an equation in y only. Then solve for y. That will let you get x.

OpenStudy (nirmalnema):

divide the 1 eq by 2 you will get x+y=5 and by arranging secong equation you will get x+y=5 so both equations are same so you will get infinitely many solutioins so option B is correct

OpenStudy (nirmalnema):

@sammy34daisy do get that

OpenStudy (anonymous):

Ahh, I see what I did wrong!